OpenStudy (anonymous):

Given , what is f(x)=-3x+4 what is f(2) ? A. –28 B. –24 C. –2 D. 10

hero (hero):

f(x) = -3(x) + 4 f(2) = -3(2) + 4

OpenStudy (anonymous):

-2?

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

Thanks y'all